• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name johna

    Log entry time 00:16:44 on April 25, 2007

    Entry number 199929

    keyword=Shift summary

    Lost about 3-4 hours due to:

    1)Access to restore missing TDC channel in left arm S1 (channel was restored and then lost again, see log entries for details)

    2)Access to reset all magnets after He flow was stopped due to compressor failure in warm return line

    3)Beam problems in Injector.

    Otherwise, data taking on 4He at 90uA. Good data runs are:

    2747-2750 (2750 is short) and 2758(short)-2764. Total of 11-12M triggers.

    PS3=2 for most runs (5% DT), set to 1 starting run 2762 (20% DT)